# A Unified, Practical, and Understandable Summary of Non-transactional Consistency Levels in Distributed Replication

### v2

@@ 6. Conclusion
This paper presents a unified, practical, and understandable summary of non-transactional consistency levels in the context of distributed data replication systems. We develop an intuitive shared object pool (SOP) model and define consistency levels within this framework by constructing them out of two types of ordering validity constraints: convergence and relationship. We explain the four most common levels, namely linearizability, sequential, causal+, and eventual consistency, along with other refined levels with detailed examples. We also discuss their availability upper bound.
As replicated, fault-tolerant object storage systems become the cloud-era norm, we believe this paper provides useful guidance for replication protocol designers and distributed system engineers.

### v4

@@ 7. Conclusion
This paper presents a unified, practical, and understandable model of non-transactional consistency levels in the context of distributed replication systems. We develop an intuitive model – the shared object pool (SOP) – and show that concise definitions of useful consistency levels can be constructed out of two types of ordering validity constraints: convergence and relationship. We demonstrate the model’s expressiveness and practical relevance with a Jepsen-integrated checker implementation. As replicated systems become the cloud-era norm, we believe this paper provides useful guidance to protocol designers and distributed system engineers.